OpenStudy (anonymous):

Stuck on a partial fraction integral problem:

OpenStudy (anonymous):

Firstly: \[\int\limits_{}^{}(x ^{2}+11)/((x-3)(x ^{2}+1)) = \int\limits_{}^{}A/(x-3) + \int\limits_{}^{}(Bx+C)/(x ^{2}+1)\] Find variables: Mulitply by LCD after setting up:\[x ^{2}+11=A(x ^{2}+1)+Bx+C(x-3)\] Distribute on right: \[= Ax ^{2}+A+Bx ^{2}-3Bx+Cx-3C\] Combine like terms: \[(A+B)x ^{2}+(-3B+C)x+(-3C+A)x ^{0}\] Equate coefficients of like terms on opposite sides of equation: \[x ^{0}: 11 = -3C+A\] \[x: 0=-3B+C\] \[x ^{2}:1=A+B\]

OpenStudy (anonymous):

Now from equating the coefficients of x I know I have C = 3B which I can then plug in for c in the top?

OpenStudy (anonymous):

The x^0 equation

OpenStudy (anonymous):

But I am not sure if that is allowed or if I am solving the system of equation it makes after plugging in

OpenStudy (anonymous):

Don't know why I can't see the latex here, so pls attach the question

OpenStudy (anonymous):

Dang it, can anyone see it?

OpenStudy (anonymous):

http://imgur.com/ZXOdQ

OpenStudy (anonymous):

There's a screen shot

OpenStudy (anonymous):

So the systems of equation I make is: 11=-9B+A <<(new version of X^0 equation, since C=3B from x equation) +9=9B+9A<<(by multiplying x^2 equation by 9) ----------- So, 19=10A : A = 19/10

OpenStudy (anonymous):

And that's where I am at right now I am skeptical I have been correct up till now so I just want to make sure before I proceed.

OpenStudy (anonymous):

I got A=2, B=-1, and C=-3

OpenStudy (anonymous):

Yeap I also got A=2,B=-1 and C=-3

OpenStudy (anonymous):

What am I doing wrong?

OpenStudy (anonymous):

Should have been Ax^2+A+Bx+Cx-3C

OpenStudy (anonymous):

All I did was distribute out during this step: A(x2+1)+Bx+C(x−3)

OpenStudy (anonymous):

A+B=1 C-3B=0 A-3C=11 let C=3B then A-3C=11 turns into A-3(3B)=11 so A-9B=11, since A+B=1 then B=1-A, so A-9B=11 turns into A-9(1-A)=11, so A-9+9A=11 so 10A=11+9, 10A=20 and A=2

OpenStudy (anonymous):

Distributing that doesn't yield: Ax^2+A+Bx+Cx-3C, that is missing the Bx^2 no?

OpenStudy (anonymous):

yes it is missing a bx^2

OpenStudy (anonymous):

your first post on the decomposition was correct\[Ax2+A+Bx2−3Bx+Cx−3C \]

OpenStudy (anonymous):

So my attempt at setting up a system by muliplying by 9 was incorrect?

OpenStudy (anonymous):

Oh wait.. it would have worked, I multiplied wrong

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Cheers

OpenStudy (anonymous):

you know how to do the rest?